# Optimal Low-Depth Quantum Signal-Processing Phase Estimation

### v1

@@ A.2 Quantum signal processing and polynomial analysis
Condition (3)”: Condition (3), which is equivalent to $\det U(x,\varOmega)=1$, directly follows the special unitarity.
Symmetric QSP”: Note that when $\Phi$ is symmetric, $U(x,\Phi)$ is invariant under the matrix transpose which reverses the order of phase factors. Using $U(x,\Phi)=U(x,\Phi)^{\top}$, the condition on the polynomial $Q(x)=Q^{*}(x)$ follows the transformation of the off-diagonal element. Therefore, $Q\in\mathbb{R}[x]$ is a real polynomial. ∎
The previous theorem bridges the gap between the periodic circuits and the analysis of polynomial. In the paper, we will frequently invoke an important inequality of polynomials to bound error, which is stated as follows.
@@
A.3 Fermionic simulation gate (FsimGate)
@@
One of the most important two-qubit quantum gates is controlled-Z gate (CZ). It forms universal gate sets with several single-qubit gates and it is a pivotal building block for demonstrating surface code [1]. CZ is in the gate class of FsimGate’s, which can be generated by setting $\theta=\varphi=\chi=\psi=0$ and $\phi=\pi$. Due to the noisy implementation of CZ, the resulting quantum gate is an FsimGate slightly deviating the perfect CZ. In order to perform high-fidelity quantum computation, one has to characterize the erroneous parameters of an FsimGate which include CZ as a special case. The characterization of gate parameters relies on quantum phase estimation techniques.
A.4 Notation
@@
Appendix B Details on QSPE in general cases
@@
Figure 8: Quantum circuit for QSPE in general cases. The input quantum state is prepared to be logical Bell states in either $\ket{+_{\ell}}$ or |$\ket{\mathrm{i}_{\ell}}$ according to the type of experiment. The quantum circuit enjoys a periodic structure of the unknown (multi-qubit) $U$-gate and a tunable logical $Z$-rotation. The measurement is performed to measure the transition probability onto the logical basis state |$\ket{0_{\ell}}$.
@@
It is worth noting that the periodic part is independent with the phase accumulation angle $\chi$. Consequently, despite the amplification of $\theta$ and $\varphi$, the angle $\chi$ is not amplified in this basis. Hence, the estimation accuracy of $\chi$ is not significantly affected by increasing the depth parameter $d$ in the quantum circuit outlined in Figure 8. To conclude, Equation 18 will be the starting point of our analysis in the later section.
Appendix C Analytical structure of periodic circuit
The QSPE circuit in Figure 1 (or Figure 8) enjoys a periodic structure by interleaving $U$-gate and $Z$-rotation. This periodic structure is studied by the theory of QSP (Theorem 3). Consequentially, the QSPE circuit admits some polynomial representation. In this section, we will derive the analytical form of the structure of the QSPE circuit. We start from the exact closed-form results of the QSPE circuit in Section C.1. In Section C.2, we derive a good approximation to the closed-form exact results. The analysis in this section proves the following theorem which summarizes findings about the structure of QSPE.
@@ Appendix C Analytical structure of periodic circuit
C.1 Exact representation of the periodic circuit
We abstract a simple $\mathrm{SU}(2)$-product model which can be shown as the building block of the QSPE circuit in Figure 1 (or Figure 8). It turns out that the model admits a polynomial representation.
@@ C.1 Exact representation of the periodic circuit
In the quantum circuit defined above, the $X$- and $Z$-rotations are interleaved, which agrees with the structure of QSP in Theorem 3. The theory of QSP implies that the $\mathrm{SU}(2)$-product model enjoys a structure representing by polynomials which is given by the following lemma.
@@
Following [10, Theorem 4], there exists two polynomials $P^{(d)}_{\omega},Q^{(d)}_{\omega}\in\mathbb{C}[x]$ so that Equation 23 holds. Because $U^{(d)}(\omega,\arccos(x))$ is a QSP unitary with a set of symmetric phase factors, $Q_{\omega}^{(d)}\in\mathbb{R}_{d-1}[x]$ is a real polynomial according to [8, Theorem 2]. Equation 24 holds by taking the determinant of Equation 23. ∎
@@
Let us prove the theorem by induction. The base case is $d=1$, where $P^{(1)}_{\omega}(x)=e^{2\mathrm{i}\omega}x$ and $Q^{(1)}_{\omega}(x)=1$. Assuming that the induction hypothesis holds for $d$, we will prove it also holds for $d+1$. Using Definitions 6 and 7, the polynomials can be determined by a recurrence relation
@@
The closed-form results above help us to analyze the dynamics of Figure 1 (or Figure 8) where we apply a Pauli $Z$ modulation $e^{\mathrm{i}\omega Z_{A_{0}}}$ to the periodic circuit. Restricted to the single-excitation subspace, the matrix representation of the QSPE circuit in Figure 1 (or Figure 8) is
@@
Using Theorems 8, 24 and 35 as intermediate steps, we have
@@
Notice that if the transition probability between tensor-product states is measured, the magnitude of the signal (the nontrivial $\theta$ dependence in the transition probability) is $\mathcal{O}\left(\sin^{2}\theta\right)$. Nonetheless, by preparing the input quantum state as Bell states, Corollary 10 reveals that the magnitude of the signal is lifted to $\mathcal{O}(\sin\theta)$ instead. Therefore, when $\theta$ is extremely small, it is a significant improvement of the SNR especially in the presence of realistic errors.
C.2 Approximate Fourier coefficients
Theorem 9 shows that the $\theta$ and $\varphi$ dependence are factored completely in the amplitude and the phase of the Fourier coefficients of the reconstructed function $\mathfrak{h}$ respectively. Given the angle $\omega$ of the $Z$-rotation is tunable, we can sample the data point by performing the QSPE circuit in Figure 1 (or Figure 8) with equally spaced angles $\omega_{j}=\frac{j}{2d-1}\pi$ where $j=0,\cdots,2d-2$. These $2(2d-1)$ quantum experiments yield two sequences of measurement probabilities $\mathbf{p}_{X}^{\mathrm{exp}}:=\left(p^{\mathrm{exp}}_{X}(\omega_{0}),p^{\mathrm{exp}}_{X}(\omega_{1}),\cdots,p^{\mathrm{exp}}_{X}(\omega_{2d-2})\right)$ and $\mathbf{p}_{Y}^{\mathrm{exp}}:=\left(p^{\mathrm{exp}}_{Y}(\omega_{0}),p^{\mathrm{exp}}_{Y}(\omega_{1}),\cdots,p^{\mathrm{exp}}_{Y}(\omega_{2d-2})\right)$. Therefore, we can compute $\mathfrak{h}^{\mathrm{exp}}=\mathbf{p}_{X}^{\mathrm{exp}}+\mathbf{p}_{Y}^{\mathrm{exp}}-\frac{1+\mathrm{i}}{2}$ from experimental data. The Fourier coefficients of $\mathfrak{h}$ can be computed by fast Fourier transform (FFT). The Fourier coefficients $\mathbf{c}^{\mathrm{exp}}:=\left(c_{-d+1}^{(\mathrm{exp})},c_{-d+2}^{(\mathrm{exp})},\cdots,c_{d-1}^{(\mathrm{exp})}\right)=\mathsf{FFT}(\mathfrak{h}^{\mathrm{exp}})$ can be computed efficiently using FFT. In order to infer $\theta$ and $\varphi$ accurately and efficiently from the data, we need to study the approximate structure of the Fourier coefficients first.
@@ C.2 Approximate Fourier coefficients
There are two implications of the previous theorem. First, it suggests that the magnitude of the Fourier coefficients of negative indices are $\mathcal{O}\left(\sin^{3}\theta\right)$. If they are included in the formalism of the inference problem in the Fourier space, the accuracy of inference may be heavily contaminated because of the nearly vanishing SNR when $\theta\ll 1$. On the other hand, the amplitude of the Fourier coefficients of nonnegative indices tightly concentrate at $\sin\theta$ when $\theta\ll 1$. Therefore, a nice linear approximation of the Fourier coefficients holds in the case of extremely small swap angle: for any $k=0,\cdots,d-1$
@@
Figure 9: Numerical justification of the approximated Fourier coefficients. We set $d=5$. The solid curve is computed by integrating the Fourier transformation. The dashed curve is derived from the approximation in the previous theorem. The shaded area stands for $d\theta\leq 1$. (a) Fourier coefficients of negative indices. (b) Fourier coefficient of nonnegative indices.
Appendix D Robust estimator against Monte Carlo sampling error
@@
Problem 12 (Calibrating $U$-gate using QSPE).
(1) QSPE: Given experimentally measured probabilities of QSPE circuits in Figure 1 (or Figure 8) on $\{\omega_{j}:j=0,\cdots,2d-2\}$, infer $\theta$ and $\varphi$ accurately.
@@
A dominant and unavoidable source of errors in quantum metrology is Monte Carlo sampling error due to the finite sample size in quantum measurements. Such limitation derives from both practical concerns of the efficiency of quantum metrology, and realistic constraints where some system parameters can drift over time and can only be monitored by sufficiently fast protocols. In this section, we analyze the effect of Monte Carlo sampling error in our proposed metrology algorithm by characterizing the sampling error as a function of quantum circuit depth, FsimGate parameters and sample size. The result will also be used in Appendix F to prove that our estimator based on QSPE is optimal. In the following analysis, we annotate with superscript “$\mathrm{exp}$” to represent experimentally measured probability as oppose to expected probability from theory.
D.1 Modeling the Monte Carlo sampling error
@@
With a characterization of Monte Carlo sampling error, we are able to measure the robustness of the signal against error by the signal-to-noise ratio (SNR). The SNR of each Fourier coefficient is defined as the ratio between the squared Fourier coefficient and the variance of its associated additive sampling error. We define the SNR of QSPE in 12 by the minimal component-wise SNR. The following theorem gives a characterization of the SNR.
@@
D.2 Statistical estimator against Monte Carlo sampling error
@@
|$\absolutevalue{\vec{c}^{\mathrm{exp}}}:=\left(\mathsf{amplitude}(c_{0}^{\mathrm{exp}}),\cdots,\mathsf{amplitude}(c_{d-1}^{\mathrm{exp}})\right)^{\top},\ \vec{\widetilde{\mymathbb{1}}}=(\underbrace{1,\cdots,1}_{d})^{\top}.$ (76)
@@
$\hat{\theta}=\mathop{\mathrm{argmin}}_{\theta}\left(\absolutevalue{\vec{c}^{\mathrm{exp}}}-\theta\vec{\widetilde{\mymathbb{1}}}\right)^{\top}\left(\mathcal{C}^{(\mathrm{amp})}\right)^{-1}\left(\absolutevalue{\vec{c}^{\mathrm{exp}}}-\theta\vec{\widetilde{\mymathbb{1}}}\right).$ (77)
@@
Remarkably, by computing the sequential phase difference, the troublesome $(2\pi)$-periodicity in Equation 74 can be overcome. According to this equation, the noise is turned to a colored noise process. Let the covariance be
@@
$\hat{\varphi}=\mathop{\mathrm{argmin}}_{\varphi}\left(\vec{\Delta}-2\varphi\vec{\mymathbb{1}}\right)^{\top}\mathcal{D}^{-1}\left(\vec{\Delta}-2\varphi\vec{\mymathbb{1}}\right)$ (80)
@@
$\vec{\Delta}=\left(\mathsf{phase}\left(c_{0}^{\mathrm{exp}}\overline{c_{1}^{\mathrm{exp}}}\right),\cdots,\mathsf{phase}\left(c_{d-2}^{\mathrm{exp}}\overline{c_{d-1}^{\mathrm{exp}}}\right)\right)^{\top},\text{ and }\vec{\mymathbb{1}}=(\underbrace{1,\cdots,1}_{d-1})^{\top}.$ (81)
@@
We first estimate the covariance of the real and imaginary components of the Monte Carlo sampling error in Fourier coefficients. Following Equations 65 and 13, for any $k\neq k^{\prime}$,
@@
Let the all-one vector be $\vec{\mymathbb{1}}=(\underbrace{1,\ldots,1}_{d-1})^{\top}$ and the discrete Laplacian matrix be
@@
$\hat{\theta}=\frac{1}{d}\sum_{k=0}^{d-1}\absolutevalue{c_{k}^{\mathrm{exp}}}\quad\text{ and }\quad\hat{\varphi}=\frac{1}{2}\frac{\vec{\mymathbb{1}}^{\top}\mathfrak{D}^{-1}\vec{\Delta}}{\vec{\mymathbb{1}}^{\top}\mathfrak{D}^{-1}\vec{\mymathbb{1}}}.$ (100)
@@
In practice, an additional moving average filter in Ref. [30] can be applied to the data to further numerically boost the SNR. For completeness, we exactly compute the optimal variance from Cramér-Rao lower bound and discuss the optimality of the estimators in Appendix F.
D.3 Improving the estimator of swap angle using the peak information provided by $\hat{\varphi}$
@@
Let $\mathfrak{D}$ be the $d$-by-$d$ discrete Laplacian matrix and $\mymathbb{1}:=(1,1,\cdots,1)\in\mathbb{R}^{d}$. The swap angle can be estimated by the statistical estimator
$\hat{\theta}_{\mathrm{pd}}=\frac{1}{2}\frac{\mymathbb{1}^{\top}\mathfrak{D}^{-1}\vec{\Gamma}}{\mymathbb{1}^{\top}\mathfrak{D}^{-1}\mymathbb{1}}.$ (103)
@@
where $\delta_{d}(\omega-\varphi,\theta)$ is the systematic error raising in the linearization of the model. Using Equations 111 and 112, when |$\absolutevalue{\omega-\varphi}\leq\frac{\pi}{d+2}$, the systematic error is bounded as
@@
$\vec{\Gamma}=\left(\Gamma_{d}^{\mathrm{exp}},\Gamma_{d+2}^{\mathrm{exp}},\cdots,\Gamma_{3d-2}^{\mathrm{exp}}\right)^{\top}\in\mathbb{R}^{d}$ (119)
@@
$\vec{\delta}(\hat{\varphi}-\varphi,\theta)=\left(\delta_{d}(\hat{\varphi}-\varphi,\theta),\delta_{d+2}(\hat{\varphi}-\varphi,\theta),\cdots,\delta_{3d-2}(\hat{\varphi}-\varphi,\theta)\right)^{\top}\in\mathbb{R}^{d}.$ (120)
@@
$\hat{\theta}_{\mathrm{pd}}=\frac{1}{2}\frac{\mymathbb{1}^{\top}\mathfrak{D}^{-1}\vec{\Gamma}}{\mymathbb{1}^{\top}\mathfrak{D}^{-1}\mymathbb{1}}.$ (121)
According to Ref. [12], the matrix-multiplication form can be exactly represented as a convex combination: for any $d$-dimensional vector $\vec{X}=(X_{0},\cdots,X_{d-1})^{\top}$
$\frac{\mymathbb{1}^{\top}\mathfrak{D}^{-1}\vec{X}}{\mymathbb{1}^{\top}\mathfrak{D}^{-1}\mymathbb{1}}=\sum_{k=0}^{d-1}\mu_{k}X_{k}$ (122)
@@
$\mathrm{Var}\left(\hat{\theta}_{\mathrm{pd}}\right)=\frac{1}{4}\frac{1}{4M}\frac{1}{\mymathbb{1}^{\top}\mathfrak{D}^{-1}\mymathbb{1}}=\frac{3}{4Md(d+1)(d+2)}\approx\frac{3}{4d^{3}M}.$ (124)
@@
$\mathbb{E}\left(\hat{\theta}_{\mathrm{pd}}\mathbb{I}_{\absolutevalue{\hat{\varphi}_{\mathrm{pri}}-\varphi}\leq\frac{\pi}{3d}}\bigg{|}\hat{\varphi}_{\mathrm{pri}}\right)=\left(\theta+\frac{1}{2}\frac{\mymathbb{1}^{\top}\mathfrak{D}^{-1}\vec{\delta}(\hat{\varphi}_{\mathrm{pri}}-\varphi,\theta)}{\mymathbb{1}^{\top}\mathfrak{D}^{-1}\mymathbb{1}}\right)\mathbb{I}_{\absolutevalue{\hat{\varphi}_{\mathrm{pri}}-\varphi}\leq\frac{\pi}{3d}}.$ (127)
@@
D.4 Peak regression and peak fitting
In order to circumvent the over-confident reliance on the a priori guess of $\varphi$, the method can be improved by regressing distinct samples with respect to analytical expressions on the unknown angle parameters. Suppose $n$ samples are made with $\{(\omega_{j},d_{j},\mathfrak{h}^{\mathrm{exp},j}):j=1,\cdots,n\}$. One can consider perform a nonlinear regression on the data to infer the unknown parameters, which is given by the following minimization problem

### v2

@@ Acknowledgments
“Condition (3)”: Condition (3), which is equivalent to $\det U(x,\varOmega)=1$, directly follows the special unitarity.
“Symmetric QSP”: Note that when $\Phi$ is symmetric, $U(x,\Phi)$ is invariant under the matrix transpose which reverses the order of phase factors. Using $U(x,\Phi)=U(x,\Phi)^{\top}$, the condition on the polynomial $Q(x)=Q^{*}(x)$ follows the transformation of the off-diagonal element. Therefore, $Q\in\mathbb{R}[x]$ is a real polynomial. ∎
The previous theorem bridges the gap between the periodic circuits and the analysis of polynomials. In the paper, we will frequently invoke an important inequality of polynomials to bound error, which is stated as follows.
@@
Supplementary Information 1.3 Fermionic simulation gate (FsimGate)
@@
One of the most important two-qubit quantum gates is controlled-Z gate (CZ). It forms universal gate sets with several single-qubit gates and it is a pivotal building block for demonstrating surface code [2]. CZ is in the gate class of FsimGate’s, which can be generated by setting $\theta=\varphi=\chi=\psi=0$ and $\phi=\pi$. Due to the noisy implementation of CZ, the resulting quantum gate is an FsimGate slightly deviating from the perfect CZ. In order to perform high-fidelity quantum computation, one has to characterize the erroneous parameters of an FsimGate which include CZ as a special case. The characterization of gate parameters relies on quantum phase estimation techniques.
Supplementary Information 1.4 Notation
@@
Supplementary Information 2 Details on QSPE in general cases
@@
Supplementary Figure 3: Quantum circuit for QSPE in general cases. The input quantum state is prepared to be logical Bell states in either $\ket{+_{\ell}}$ or |$\ket{\mathrm{i}_{\ell}}$ according to the type of experiment. The quantum circuit enjoys a periodic structure of the unknown (multi-qubit) $U$-gate and a tunable logical $Z$-rotation. The measurement is performed to measure the transition probability onto the logical basis state |$\ket{0_{\ell}}$.
@@
It is worth noting that the periodic part is independent of the phase accumulation angle $\chi$. Consequently, despite the amplification of $\theta$ and $\varphi$, the angle $\chi$ is not amplified on this basis. Hence, the estimation accuracy of $\chi$ is not significantly affected by increasing the depth parameter $d$ in the quantum circuit outlined in Figure 3. To conclude, Equation 6 will be the starting point of our analysis in the later section.
Supplementary Information 3 Analytical structure of periodic circuit
The QSPE circuit in Figure 1 in the main text (or Figure 3) enjoys a periodic structure by interleaving $U$-gate and $Z$-rotation. This periodic structure is studied by the theory of QSP (Theorem 1). Consequentially, the QSPE circuit admits some polynomial representation. In this section, we will derive the analytical form of the structure of the QSPE circuit. We start from the exact closed-form results of the QSPE circuit in Section Supplementary Information 3.1. In Section Supplementary Information 3.2, we derive a good approximation to the closed-form exact results. The analysis in this section proves the following theorem which summarizes findings about the structure of QSPE.
@@
Supplementary Information 3.1 Exact representation of the periodic circuit
We abstract a simple $\mathrm{SU}(2)$-product model which can be shown as the building block of the QSPE circuit in Figure 1 in the main text (or Figure 3). It turns out that the model admits a polynomial representation.
@@
In the quantum circuit defined above, the $X$- and $Z$-rotations are interleaved, which agrees with the structure of QSP in Theorem 1. The theory of QSP implies that the $\mathrm{SU}(2)$-product model enjoys a structure represented by polynomials which is given by the following lemma.
@@
Following [11, Theorem 4], there exists two polynomials $P^{(d)}_{\omega},Q^{(d)}_{\omega}\in\mathbb{C}[x]$ so that Equation 11 holds. Because $U^{(d)}(\omega,\arccos(x))$ is a QSP unitary with a set of symmetric phase factors, $Q_{\omega}^{(d)}\in\mathbb{R}_{d-1}[x]$ is a real polynomial according to [9, Theorem 2]. Equation 12 holds by taking the determinant of Equation 11. ∎
@@
Let us prove the theorem by induction. The base case is $d=1$, where $P^{(1)}_{\omega}(x)=e^{2\mathrm{i}\omega}x$ and $Q^{(1)}_{\omega}(x)=1$. Assuming that the induction hypothesis holds for $d$, we will prove it also holds for $d+1$. Using Definitions 4 and 5, the polynomials can be determined by a recurrence relation
@@
The closed-form results above help us to analyze the dynamics of Figure 1 in the main text (or Figure 3) where we apply a Pauli $Z$ modulation $e^{\mathrm{i}\omega Z_{A_{0}}}$ to the periodic circuit. Restricted to the single-excitation subspace, the matrix representation of the QSPE circuit in Figure 1 in the main text (or Figure 3) is
@@
Using Theorems 6, 12 and 23 as intermediate steps, we have
@@
Notice that if the transition probability between tensor-product states is measured, the magnitude of the signal (the nontrivial $\theta$ dependence in the transition probability) is $\mathcal{O}\left(\sin^{2}\theta\right)$. Nonetheless, by preparing the input quantum state as Bell states, Corollary 8 reveals that the magnitude of the signal is lifted to $\mathcal{O}(\sin\theta)$ instead. Therefore, when $\theta$ is small, it is a significant improvement of the SNR especially in the presence of realistic errors.
Supplementary Information 3.2 Approximate Fourier coefficients
Theorem 7 shows that the $\theta$ and $\varphi$ dependence are factored completely in the amplitude and the phase of the Fourier coefficients of the reconstructed function $\mathfrak{h}$ respectively. Given the angle $\omega$ of the $Z$-rotation is tunable, we can sample the data point by performing the QSPE circuit in Figure 1 in the main text (or Figure 3) with equally spaced angles $\omega_{j}=\frac{j}{2d-1}\pi$ where $j=0,\cdots,2d-2$. These $2(2d-1)$ quantum experiments yield two sequences of measurement probabilities $\mathbf{p}_{X}^{\mathrm{exp}}:=\left(p^{\mathrm{exp}}_{X}(\omega_{0}),p^{\mathrm{exp}}_{X}(\omega_{1}),\cdots,p^{\mathrm{exp}}_{X}(\omega_{2d-2})\right)$ and $\mathbf{p}_{Y}^{\mathrm{exp}}:=\left(p^{\mathrm{exp}}_{Y}(\omega_{0}),p^{\mathrm{exp}}_{Y}(\omega_{1}),\cdots,p^{\mathrm{exp}}_{Y}(\omega_{2d-2})\right)$. Therefore, we can compute $\mathfrak{h}^{\mathrm{exp}}=\mathbf{p}_{X}^{\mathrm{exp}}+\mathbf{p}_{Y}^{\mathrm{exp}}-\frac{1+\mathrm{i}}{2}$ from experimental data. The Fourier coefficients of $\mathfrak{h}$ can be computed by fast Fourier transform (FFT). The Fourier coefficients $\mathbf{c}^{\mathrm{exp}}:=\left(c_{-d+1}^{(\mathrm{exp})},c_{-d+2}^{(\mathrm{exp})},\cdots,c_{d-1}^{(\mathrm{exp})}\right)=\mathsf{FFT}(\mathfrak{h}^{\mathrm{exp}})$ can be computed efficiently using FFT. In order to infer $\theta$ and $\varphi$ accurately and efficiently from the data, we need to study the approximate structure of the Fourier coefficients first.
@@
There are two implications of the previous theorem. First, it suggests that the magnitude of the Fourier coefficients of negative indices are $\mathcal{O}\left(\sin^{3}\theta\right)$. If they are included in the formalism of the inference problem in the Fourier space, the accuracy of inference may be heavily contaminated because of the nearly vanishing SNR when $\theta\ll 1$. On the other hand, the amplitude of the Fourier coefficients of nonnegative indices tightly concentrates at $\sin\theta$ when $\theta\ll 1$. Therefore, a nice linear approximation of the Fourier coefficients holds in the case of small swap angle: for any $k=0,\cdots,d-1$
@@
Supplementary Figure 4: Numerical justification of the approximated Fourier coefficients. We set $d=5$. The solid curve is computed by integrating the Fourier transformation. The dashed curve is derived from the approximation in the previous theorem. The shaded area stands for $d\theta\leq 1$. (a) Fourier coefficients of negative indices. (b) Fourier coefficient of nonnegative indices.
Supplementary Information 4 Robust estimator against Monte Carlo sampling error
@@
Problem 10 (Calibrating $U$-gate using QSPE).
(1) QSPE: Given experimentally measured probabilities of QSPE circuits in Figure 1 in the main text (or Figure 3) on $\{\omega_{j}:j=0,\cdots,2d-2\}$, infer $\theta$ and $\varphi$ accurately.
@@
A dominant and unavoidable source of errors in quantum metrology is Monte Carlo sampling error due to the finite sample size in quantum measurements. Such limitation derives from both practical concerns of the efficiency of quantum metrology, and realistic constraints where some system parameters can drift over time and can only be monitored by sufficiently fast protocols. In this section, we analyze the effect of Monte Carlo sampling error in our proposed metrology algorithm by characterizing the sampling error as a function of quantum circuit depth, FsimGate parameters and sample size. The result will also be used in Section Supplementary Information 6 to prove that our estimator based on QSPE is optimal. In the following analysis, we annotate with superscript “$\mathrm{exp}$” to represent experimentally measured probability as opposed to expected probability from theory.
Supplementary Information 4.1 Modeling the Monte Carlo sampling error
@@
With a characterization of Monte Carlo sampling error, we are able to measure the robustness of the signal against error by the signal-to-noise ratio (SNR). The SNR of each Fourier coefficient is defined as the ratio between the squared Fourier coefficient and the variance of its associated additive sampling error. We define the SNR of QSPE in 10 by the minimal component-wise SNR. The following theorem gives a characterization of the SNR.
@@
Supplementary Information 4.2 Statistical estimator against Monte Carlo sampling error
@@
|$\absolutevalue{\mathbf{c}^{\mathrm{exp}}}:=\left(\mathsf{amplitude}(c_{0}^{\mathrm{exp}}),\cdots,\mathsf{amplitude}(c_{d-1}^{\mathrm{exp}})\right)^{\top},\ \bm{\widetilde{\bm{1}}}=(\underbrace{1,\cdots,1}_{d})^{\top}.$ (64)
@@
$\hat{\theta}=\mathop{\mathrm{argmin}}_{\theta}\left(\absolutevalue{\mathbf{c}^{\mathrm{exp}}}-\theta\bm{\widetilde{\bm{1}}}\right)^{\top}\left(\mathcal{C}^{(\mathrm{amp})}\right)^{-1}\left(\absolutevalue{\mathbf{c}^{\mathrm{exp}}}-\theta\bm{\widetilde{\bm{1}}}\right).$ (65)
@@
Remarkably, by computing the sequential phase difference, the troublesome $(2\pi)$-periodicity in Equation 62 can be overcome. According to this equation, the noise is turned into a colored noise process. Let the covariance be
@@
$\hat{\varphi}=\mathop{\mathrm{argmin}}_{\varphi}\left(\bm{\Delta}-2\varphi\bm{\bm{1}}\right)^{\top}\mathcal{D}^{-1}\left(\bm{\Delta}-2\varphi\bm{\bm{1}}\right)$ (68)
@@
$\bm{\Delta}=\left(\mathsf{phase}\left(c_{0}^{\mathrm{exp}}\overline{c_{1}^{\mathrm{exp}}}\right),\cdots,\mathsf{phase}\left(c_{d-2}^{\mathrm{exp}}\overline{c_{d-1}^{\mathrm{exp}}}\right)\right)^{\top},\text{ and }\bm{\bm{1}}=(\underbrace{1,\cdots,1}_{d-1})^{\top}.$ (69)
@@
We first estimate the covariance of the real and imaginary components of the Monte Carlo sampling error in Fourier coefficients. Following Equations 53 and 11, for any $k\neq k^{\prime}$,
@@
Let the all-one vector be $\bm{\bm{1}}=(\underbrace{1,\ldots,1}_{d-1})^{\top}$ and the discrete Laplacian matrix be
@@
$\hat{\theta}=\frac{1}{d}\sum_{k=0}^{d-1}\absolutevalue{c_{k}^{\mathrm{exp}}}\quad\text{ and }\quad\hat{\varphi}=\frac{1}{2}\frac{\bm{\bm{1}}^{\top}\mathfrak{D}^{-1}\bm{\Delta}}{\bm{\bm{1}}^{\top}\mathfrak{D}^{-1}\bm{\bm{1}}}.$ (88)
@@
In practice, an additional moving average filter in Ref. [33] can be applied to the data to further numerically boost the SNR. For completeness, we exactly compute the optimal variance from Cramér-Rao lower bound and discuss the optimality of the estimators in Section Supplementary Information 6.
Supplementary Information 4.3 Improving the estimator of swap angle using the peak information provided by $\hat{\varphi}$
@@
Let $\mathfrak{D}$ be the $d$-by-$d$ discrete Laplacian matrix and $\bm{1}:=(1,1,\cdots,1)\in\mathbb{R}^{d}$. The swap angle can be estimated by the statistical estimator
$\hat{\theta}_{\mathrm{pd}}=\frac{1}{2}\frac{\bm{1}^{\top}\mathfrak{D}^{-1}\bm{\Gamma}}{\bm{1}^{\top}\mathfrak{D}^{-1}\bm{1}}.$ (91)
@@
where $\delta_{d}(\omega-\varphi,\theta)$ is the systematic error raising in the linearization of the model. Using Equations 99 and 100, when |$\absolutevalue{\omega-\varphi}\leq\frac{\pi}{d+2}$, the systematic error is bounded as
@@
$\bm{\Gamma}=\left(\Gamma_{d}^{\mathrm{exp}},\Gamma_{d+2}^{\mathrm{exp}},\cdots,\Gamma_{3d-2}^{\mathrm{exp}}\right)^{\top}\in\mathbb{R}^{d}$ (107)
@@
$\mathrm{\bm{\delta}}(\hat{\varphi}-\varphi,\theta)=\left(\delta_{d}(\hat{\varphi}-\varphi,\theta),\delta_{d+2}(\hat{\varphi}-\varphi,\theta),\cdots,\delta_{3d-2}(\hat{\varphi}-\varphi,\theta)\right)^{\top}\in\mathbb{R}^{d}.$ (108)
@@
$\hat{\theta}_{\mathrm{pd}}=\frac{1}{2}\frac{\bm{1}^{\top}\mathfrak{D}^{-1}\bm{\Gamma}}{\bm{1}^{\top}\mathfrak{D}^{-1}\bm{1}}.$ (109)
According to Ref. [14], the matrix-multiplication form can be exactly represented as a convex combination: for any $d$-dimensional vector $\mathbf{X}=(X_{0},\cdots,X_{d-1})^{\top}$
$\frac{\bm{1}^{\top}\mathfrak{D}^{-1}\mathbf{X}}{\bm{1}^{\top}\mathfrak{D}^{-1}\bm{1}}=\sum_{k=0}^{d-1}\mu_{k}X_{k}$ (110)
@@
$\mathrm{Var}\left(\hat{\theta}_{\mathrm{pd}}\right)=\frac{1}{4}\frac{1}{4M}\frac{1}{\bm{1}^{\top}\mathfrak{D}^{-1}\bm{1}}=\frac{3}{4Md(d+1)(d+2)}\approx\frac{3}{4d^{3}M}.$ (112)
@@
$\mathbb{E}\left(\hat{\theta}_{\mathrm{pd}}\mathbb{I}_{\absolutevalue{\hat{\varphi}_{\mathrm{pri}}-\varphi}\leq\frac{\pi}{3d}}\bigg{|}\hat{\varphi}_{\mathrm{pri}}\right)=\left(\theta+\frac{1}{2}\frac{\bm{1}^{\top}\mathfrak{D}^{-1}\bm{\delta}(\hat{\varphi}_{\mathrm{pri}}-\varphi,\theta)}{\bm{1}^{\top}\mathfrak{D}^{-1}\bm{1}}\right)\mathbb{I}_{\absolutevalue{\hat{\varphi}_{\mathrm{pri}}-\varphi}\leq\frac{\pi}{3d}}.$ (115)
@@
Supplementary Information 4.4 Peak regression and peak fitting
In order to circumvent the over-confident reliance on the a priori guess of $\varphi$, the method can be improved by regressing distinct samples with respect to analytical expressions on the unknown angle parameters. Suppose $n$ samples are made with $\{(\omega_{j},d_{j},\mathfrak{h}^{\mathrm{exp},j}):j=1,\cdots,n\}$. One can consider performing a nonlinear regression on the data to infer the unknown parameters, which is given by the following minimization problem